OpenStudy (anonymous):

How can you choose resistors to decrease total voltage to one quarter of the originial voltage value?

OpenStudy (anonymous):

low resistance resistors

OpenStudy (anonymous):

I=V/R So if we want have 1/4 voltage just divided R on 1/4 or *4 :) 1/4I = V/4R (Multiplication R to 4)